# Knowledge Integration Strategies in Autonomous Vehicle Prediction and Planning: A Comprehensive Survey

### v1

@@ II-C Large Language Model and Retrieval-Augmented Generation Methods
Large language models (LLMs) and retrieval-augmented generation (RAG) techniques have emerged as transformative tools in autonomous driving. They use unstructured text and regulatory data to improve decision-making, adaptability, and interpretability. These methods enable systems to incorporate vast knowledge repositories dynamically, generate human-readable explanations, and adapt to region-specific traffic laws, addressing key challenges in prediction and planning.
A prominent application of LLMs is their ability to interpret and adapt to traffic regulations in real-time. LLaDA (Large Language Driving Assistant) [24] exemplifies this capability, using LLMs to adapt driving behavior to new environments, customs, and laws. By interpreting traffic rules from local driver handbooks, LLaDA provides context-aware driving instructions and demonstrates zero-shot generalization. Similarly, [25] propose a framework where a Traffic Regulation Retrieval Agent employs RAG to query regional traffic documents, and an LLM-based reasoning module evaluates maneuvers for compliance and safety. While this approach dynamically adapts to region-specific laws and generates rule-referenced explanations, it faces latency and LLM hallucination challenges. The TR2MTL framework [26] further advances this paradigm by using LLMs in a chain-of-thought learning paradigm to translate natural language traffic rules into metric temporal logic formulas, reducing manual rule specification and enabling scalable regulatory updates.
@@
Recent works explore using LLMs as central decision-making components in autonomous driving. DriveLLM [30] charts a comprehensive roadmap for integrating LLMs into fully autonomous driving systems, emphasizing their role in high-level reasoning, planning, and interaction with dynamic environments. By leveraging LLMs’ generative and contextual understanding capabilities, DriveLLM demonstrates how these models can enhance adaptability and interpretability in complex driving scenarios. Similarly, LanguageMPC [31] proposes a framework where LLMs serve as decision-makers within a model predictive control (MPC) pipeline. This approach combines the reasoning capabilities of LLMs with the robustness of traditional control methods, enabling the system to generate human-like driving strategies while ensuring safety and compliance. Both frameworks highlight the potential of LLMs to bridge the gap between high-level reasoning and low-level control. However, challenges such as real-time performance and integration with existing systems remain.
@@ II-D Formal Logic-Based Methods
Formal logic-based methods have become a cornerstone in autonomous driving research, offering rigorous frameworks for monitoring, verification, prediction, and planning through mathematically grounded representations of traffic rules and safety requirements. A central theme in this work is the formalization of traffic regulations using temporal logic and answer set programming (ASP). For instance, [35] formalizes traffic rules at uncontrolled intersections by encoding yielding and right–of–way as temporal logic formulas, ensuring that vehicle behaviors adhere to safe bounds even though extending this approach to complex urban scenarios remains challenging. Similarly, AUTO-DISCERN [36] employs ASP via the goal-directed s(CASP) system to simulate commonsense reasoning in driving decisions. By encoding driving rules as explicit logical constraints and generating proof trees to justify maneuvers, AUTO-DISCERN achieves high interpretability while grappling with scalability issues and sensor noise.
Another emerging direction involves integrating formal logic with data-driven techniques to enhance robustness in trajectory prediction. RuleFuser [37] exemplifies this trend by combining neural predictors with classical rule-based predictors; here, formal logic is used to represent traffic rules that guide predictions in out–of–distribution scenarios. In a related effort, the TR2MTL framework [26] leverages large language models (LLMs) to automatically translate unstructured natural language traffic rules into metric temporal logic (MTL) formulas. By employing a chain–of–thought in–context learning strategy, TR2MTL reduces manual rule formalization, although its performance is sensitive to prompt design and input quality.
Formal logic is also employed to enhance situational awareness and decision–making. [38] apply ASP–based reasoning to model and assess driver situation awareness in semi-autonomous vehicles, formalizing scene interpretation as logical rules that verify correct perception and reaction to dynamic traffic conditions. Complementing these approaches works such as [39, 40] integrate set–based reachability analysis with convex optimization to plan safe trajectories that comply with legal driving standards, albeit with challenges in accurately modeling the behavior of other road users in unpredictable environments.
Further reinforcing the role of formal methods, safety verification techniques based on differential dynamic logic [41] have been applied to prove the safety of vehicle control systems under varied driving conditions. Such methods, along with survey works like [23], illustrate that while formal logic-based approaches provide strong guarantees and a high level of interpretability, they often face challenges in terms of computational complexity, scalability, and real-time integration of continuous changing sensor inputs.
@@
In practice, many of these formal logic-based methods are integrated into hybrid and optimization–based models. Formal logic offers a versatile toolkit for ensuring provably safe and compliant autonomous driving behaviors, whether employed as constraints in loss functions, integrated into reinforcement learning reward structures, or used in optimization layers for planning. Despite their promise, ongoing research continues to address the trade-offs between the rigor of formal methods and the need for real-time, scalable performance in dynamic driving environments.
II-D1 Formal Logic in MPC and Optimization
@@
Model Predictive Control (MPC) is a widely adopted optimization-based control strategy in autonomous driving. It solves an optimal control problem over a finite receding horizon to generate trajectories that account for both current sensor data and predicted future states. This approach translates traffic rules into mathematical inequalities or cost functions directly embedded into optimization-based planning frameworks. By formulating safety rules as constraints within MPC, recent research has increasingly focused on integrating formalized traffic rules into these optimization frameworks to enhance both performance and safety. Embedding traffic constraints as mathematical inequalities and temporal logic conditions allows MPC-based planners to enforce rigorous safety and regulatory requirements while optimizing for efficiency.
One line of work translates human–interpretable traffic rules into machine–readable specifications. For example, [44] proposes a methodology that converts natural language traffic rules into Linear Temporal Logic (LTL) specifications. This translation facilitates systematic testing and verification of autonomous behaviors and provides a formal basis for integrating these rules directly into planning algorithms. In parallel, [45] develop control frameworks for traffic networks where signalized intersections are managed by synthesizing control policies that satisfy temporal logic constraints. Their approach leverages formal methods and model checking to generate strategies that guarantee compliance with complex operational constraints, thereby improving safety and efficiency in urban traffic management.
Another influential approach merges MPC with Signal Temporal Logic (STL) specifications. In [46], STL constraints are encoded as mixed–integer linear constraints within an MPC framework. This formulation allows the controller to systematically enforce temporal properties during receding horizon control, which is particularly beneficial for applications requiring stringent timing and safety adherence. Extending these ideas to the urban context, [47] apply MPC to optimize traffic signal operations under temporal logic constraints, ensuring that predefined temporal conditions are met to reduce congestion and improve traffic flow. Further advancing robustness, [48] integrates predictive robustness into STL–based MPC formulations, providing enhanced reliability against uncertainties in dynamic environments.
@@
Collectively, these studies illustrate a significant trend: the integration of formal traffic rules into optimization–based control strategies yields control systems that are both flexible and provably safe. The combined use of MPC and temporal logic—whether through LTL or STL—offers a promising pathway toward designing controllers capable of real-time adaptation while rigorously satisfying safety and regulatory requirements. This hybridization of formal logic and optimization advances autonomous driving and paves the way for future developments in intelligent transportation systems.
@@ II-E Hybrid Approaches Combining Symbolic and Neural Methods
Hybrid approaches for autonomous driving have evolved by fusing the adaptability of neural networks with the rigor and interpretability of symbolic reasoning. Such methods aim to leverage the pattern recognition power of deep learning while enforcing explicit, rule-based constraints, ultimately improving safety, robustness, and explainability. These approaches can be broadly grouped by their integration strategy.
@@
One prominent strand couples deep neural perception pipelines with symbolic reasoning modules. For example, AUTO–DISCERN [36] processes sensory inputs through deep networks whose features feed into an answer set programming–based module (via the goal-directed s(CASP) system) to apply commonsense rules and enforce driving behaviors. Similarly, the framework in [61] combines explicitly encoded traffic rules with learning–based trajectory predictors; here, the symbolic component acts as a constraint, guiding the neural network towards rule–compliant predictions while reducing the data requirements.
@@
A second category focuses on integrating symbolic knowledge directly into the learning process. In [49], symbolic constraints are injected into the loss functions and network architectures, enhancing robustness in rare or hazardous scenarios. This idea is further developed in [50], where formal rules expressed as temporal logic is incorporated in a differentiable manner, penalizing non–compliant predictions. In a related effort, [58] employs signal temporal logic and syntax trees as features within a generative adversarial network framework to improve trajectory prediction accuracy without biasing the system away from lawful behavior.
@@
A recent trend leverages large language models (LLMs) to bridge high–level symbolic reasoning and sensor-driven perception. For instance, [32] utilizes an LLM to generate high–level symbolic descriptions of the driving context, which are then fused with deep learning–based sensor interpretations. Building on this concept, DILU [51] automatically generates symbolic rules from unstructured data via LLMs and integrates them into the decision–making pipeline. Although promising, these systems still face challenges such as LLM hallucinations, integration latency, and the synchronization of subsymbolic and symbolic processing streams.
@@
Complementing these technical approaches, several surveys and alternative frameworks provide a broader perspective on the hybrid paradigm. Reviews in [52] and the survey on hybrid motion planning [53] underscore the potential of embedding domain knowledge into learning–based systems to enhance safety and interpretability. Additional work such as [40] employs set–based prediction methods—using formalized traffic rules and nondeterministic motion models—to perform reachability analysis for collision avoidance, while neural–symbolic methods like Logic Tensor Networks [54] and DeepProbLog [55] offer frameworks for integrating logical constraints into deep learning models.
@@
Finally, specialized systems have been designed to address real-time adaptability and conflict resolution. For example, [56] formalizes explicit traffic rules to rule out infeasible maneuvers, and [57] combines incremental symbolic perception with reactive control synthesis to generate verified, rule–compliant driving maneuvers in dynamic environments.
@@
In summary, hybrid approaches that integrate symbolic and neural methods offer significant advantages by combining explicit rule-based reasoning with the flexibility of deep learning. While these methods enhance interpretability and safety in complex driving scenarios, challenges remain in managing computational complexity, ensuring real-time performance, and effectively synchronizing heterogeneous components.
@@
II-E1 Diffusion and Other Learning-Based Prediction Incorporating Traffic Rules
@@
Since we focus on representing traffic rules and integrating them into learning-based methods, our approach aligns closely with hybrid methods that blend formal constraints with data-driven models.
@@ III Discussion and Future Directions
A central focus lies in emerging foundation model- and diffusion model-based approaches, which leverage unstructured data and generative mechanisms to encode traffic dynamics or explicitly integrate domain knowledge implicitly. These methods show promise in enhancing prediction accuracy and planning robustness in complex scenarios, signaling a paradigm shift in autonomous systems. Our review equips researchers with methodologies to address prediction-planning challenges while balancing safety, efficiency, and goal achievement.
Critical challenges persist. Future work must resolve real-time performance in dynamic environments, improve the interpretability of deep learning systems, and unify heterogeneous data streams (e.g., sensor inputs and traffic rules). The scalability and adaptability of foundation/diffusion models in diverse driving conditions require further validation. Multimodal reasoning and formal safety guarantees—particularly for handling nuanced, conflicting rules and hierarchical regulatory frameworks—will grow in importance.
In summary, by synthesizing state-of-the-art methods across different paradigms, this survey highlights the current capabilities and limitations in trajectory prediction and planning. It lays out a roadmap for future investigations. Advancements in integrating explicit knowledge—whether through symbolic reasoning, hybrid models, or generative approaches—will be crucial in developing the next generation of autonomous driving systems that are both safe and efficient in real-world operations. By diligently advancing knowledge representation techniques and fostering synergistic combinations of different approaches, we can pave the way for more intelligent, trustworthy, and ultimately safer autonomous vehicles that can seamlessly and reliably navigate the full spectrum of real-world driving complexities.
IV Acknowledgments
This work has been partially funded by the German Federal Ministry for Economic Affairs and Climate Action within the project “nxtAIM”. We sincerely thank Dr. Stefan Zwicklbauer for his constant support and mentorship throughout the work.

### v2

@@ II-C Large Language Model and Retrieval-Augmented Generation Methods
LLMs and RAG techniques have become valuable tools for enhancing prediction and planning systems. By automating the transformation of unstructured and regulatory data into intermediate representations, they improve decision-making, adaptability, and interpretability. These approaches enable systems to access vast knowledge bases dynamically, generate human-readable explanations, and adapt to region-specific traffic regulations, effectively addressing critical challenges in predictive and planning tasks.
A prominent application of LLMs is their ability to interpret and adapt to traffic regulations in real time. LLaDA (Large Language Driving Assistant) [29] exemplifies this capability, using LLMs to adapt driving behavior to new environments, customs, and laws. By interpreting traffic rules from local driver handbooks, LLaDA provides context-aware driving instructions and demonstrates zero-shot generalization. Similarly, [30] propose a framework where a Traffic Regulation Retrieval Agent employs RAG to query regional traffic documents and an LLM-based reasoning module evaluates maneuvers for compliance and safety. While this approach dynamically adapts to region-specific laws and generates rule-referenced explanations, it faces latency and LLM hallucination challenges. The TR2MTL framework [31] further advances this paradigm by using LLMs in a chain-of-thought learning paradigm to translate natural language traffic rules into metric temporal logic formulas, reducing manual rule specification and enabling scalable regulatory updates.
@@
Recent works explore using LLMs as central decision-making components in AD. DriveLLM [35] charts a comprehensive roadmap for integrating LLMs into fully autonomous driving systems, emphasizing their role in high-level reasoning, planning, and interaction, while the DiLu framework [40] specifically leverages LLM reasoning and reflection modules for knowledge-driven decision-making using commonsense and accumulated experience. Similarly, LanguageMPC [36] proposes a framework where LLMs serve as decision-makers within a model predictive control (MPC) pipeline. This approach combines the reasoning capabilities of LLMs with the robustness of traditional control methods, enabling the system to generate human-like driving strategies while ensuring safety and compliance. Both frameworks highlight the potential of LLMs to bridge the gap between high-level reasoning and low-level control. However, challenges such as real-time performance and integration with existing systems remain.
@@ II-D Formal Logic-Based Methods
Formal logic-based methods have become a cornerstone in autonomous driving research, offering rigorous frameworks for monitoring, verification, prediction, and planning through mathematically grounded representations of traffic rules and safety requirements. A central theme in this work is the formalization of traffic regulations using temporal logic and answer set programming (ASP). For instance, [41] formalizes traffic rules at uncontrolled intersections by encoding yielding and right–of–way as temporal logic formulas, ensuring that vehicle behaviors adhere to safe bounds, even though extending this approach to complex urban scenarios remains challenging and [49] formalized traffic regulations for dual carriageways in linear temporal logic (LTL), providing a machine-interpretable rule set. Similarly, AUTO-DISCERN [42] employs ASP via the goal-directed s(CASP) system to simulate commonsense reasoning in driving decisions. By encoding driving rules as explicit logical constraints and generating proof trees to justify maneuvers, AUTO-DISCERN achieves high interpretability while grappling with scalability issues and sensor noise.
Another emerging direction involves integrating formal logic with data-driven techniques to enhance robustness in trajectory prediction. RuleFuser [65] exemplifies this trend by combining neural predictors with classical rule-based predictors; here, formal logic is used to represent traffic rules that guide predictions in out–of–distribution scenarios.
Formal logic is also employed to enhance situational awareness and decision-making. [43] applies ASP–based reasoning to model and assess driver situation awareness in semi-autonomous vehicles, formalizing scene interpretation as logical rules that verify correct perception and reaction to dynamic traffic conditions. Further reinforcing the role of formal methods, safety verification techniques based on differential dynamic logic [46] have been applied to prove the safety of vehicle control systems under varied driving conditions. Such methods, along with survey works like [4], illustrate that while formal logic-based approaches provide strong guarantees and a high level of interpretability, they often face challenges in terms of computational complexity, scalability, and real-time integration of continuously changing sensor inputs.
@@
In practice, many of these formal logic-based methods are integrated into hybrid and optimization–based models. Formal logic offers a versatile toolkit for ensuring provably safe and compliant autonomous driving behaviors, whether employed as constraints in loss functions, integrated into reinforcement learning reward structures, or used in optimization layers for planning. However, manual formalization is a tedious process. To solve this problem, TR2MTL [31] attempts to automate this process using LLM and treating text to formal logic as a translation task. Despite their promise, ongoing research continues to address the trade-offs between the rigor of formal methods and the need for real-time, scalable performance in dynamic driving environments.
@@
Formal Logic in MPC and Optimization. Model Predictive Control (MPC) is a widely adopted optimization-based control strategy in autonomous driving. It solves an optimal control problem over a finite receding horizon to generate trajectories that account for both current sensor data and predicted future states. This approach translates traffic rules into mathematical inequalities or cost functions directly embedded into optimization-based planning frameworks. By formulating safety rules as constraints within MPC, recent research has increasingly focused on integrating formalized traffic rules into these optimization frameworks to enhance both performance and safety. Embedding traffic constraints as mathematical inequalities and temporal logic conditions allows MPC-based planners to enforce rigorous safety and regulatory requirements while optimizing for efficiency. Another influential approach merges MPC with Signal Temporal Logic (STL) specifications. In [51], STL constraints are encoded as mixed-integer linear constraints within an MPC framework. This formulation allows the controller to systematically enforce temporal properties during receding horizon control, which is particularly beneficial for applications requiring stringent timing and safety adherence. Extending these ideas to the urban context, [52] applies MPC to optimize traffic signal operations under temporal logic constraints, ensuring that predefined temporal conditions are met to reduce congestion and improve traffic flow. Building on the concept of robustness, [53] incorporates predictive robustness within an STL-based MPC framework, enhancing the system’s reliability in the face of uncertainties in dynamic environments.
Set‐Based Reachability Methods compute over‐approximations of all possible future states (reachable sets)—to guarantee collision avoidance across all admissible maneuvers. [44] combines convex over‐approximations of the ego vehicle’s reachable set with real‐time convex optimization, yielding faster planning as scenarios are constrained. [45] extends this concept to prediction, formally enclosing visible and occluded agent behaviors via nondeterministic motion models and traffic‐rule constraints. These approaches complement temporal-logic MPC planners by providing provable safety envelopes over entire state tubes rather than single nominal trajectories.
@@
Control frameworks for traffic networks, where signalized intersections are managed by synthesizing control policies that satisfy temporal logic constraints, are developed by [50]. Their approach leverages formal methods and model checking to generate strategies that guarantee compliance with complex operational constraints, thereby improving safety and efficiency in urban traffic management. Collectively, these studies illustrate a significant trend: the integration of formal traffic rules into optimization–based control strategies yields control systems that are both flexible and provably safe. The combined use of MPC and temporal logic—whether through LTL or STL—offers a promising pathway toward designing controllers capable of real-time adaptation while rigorously satisfying safety and regulatory requirements. This hybridization of formal logic and optimization advances autonomous driving and paves the way for future developments in intelligent transportation systems. For a comprehensive survey explicitly focused on formal methods for traffic rule compliance in autonomous driving, readers are referred to [4].
@@ II-E Hybrid Approaches Combining Symbolic and Neural Methods
Hybrid approaches for autonomous driving have evolved by combining the adaptability of neural networks with the rigor and interpretability of symbolic reasoning. These methods leverage the pattern recognition power of deep learning while enforcing implicit and explicit rule-based constraints, enhancing safety, robustness, and explainability. These approaches can be broadly categorized by their integration strategy.
@@
One notable method, AUTO-DISCERN [42], integrates deep neural perception pipelines with symbolic reasoning modules. They process sensor inputs through deep networks, whose features feed into an answer set programming-based module (using the goal-directed s(CASP) system) to apply commonsense rules and regulate driving behaviors.
@@
A second category focuses on integrating symbolic knowledge directly into the learning process. In [54], symbolic constraints are injected into the loss functions and network architectures, enhancing robustness in rare or hazardous scenarios. This idea is further developed in [55], where formal rules expressed as temporal logic is incorporated in a differentiable manner, penalizing non–compliant predictions. In a related effort, [62] employs signal temporal logic and syntax trees as features within a generative adversarial network framework to improve trajectory prediction accuracy without biasing the system away from lawful behavior. Similarly, the SHIFT [60] combines explicitly encoded traffic rules with learning–based uncertainty-aware trajectory predictors; here, traffic rules are used as a soft constraint and provide prior information to the predictor, guiding the neural network towards rule–compliant predictions while reducing the data requirements.
Complementing methods that inject explicit symbolic rules, another form of hybrid integration involves learning implicit behavioral rules or preferences directly from data. For example, approaches using Inverse Reinforcement Learning (IRL) infer complex reward functions from human driving demonstrations, capturing nuanced objectives that may implicitly include safety and comfort considerations, effectively learning the ‘knowledge’ embedded in expert behavior [63].
@@
A recent trend leverages LLMs to bridge high–level symbolic reasoning and sensor-driven perception. For instance, [37] utilizes an LLM to generate high–level symbolic descriptions of the driving context, which are then fused with deep learning–based sensor interpretations. Building on this concept, DILU [40] automatically generates symbolic rules from unstructured data via LLMs and integrates them into the decision–making pipeline. Although promising, these systems still face challenges such as LLM hallucinations, integration latency, and the synchronization of subsymbolic and symbolic processing streams.
@@
Complementing these technical approaches, several surveys and alternative frameworks provide a broader perspective on the hybrid paradigm. Reviews in [56] and the survey on hybrid motion planning [57] underscore the potential of embedding domain knowledge into learning–based systems to enhance safety and interpretability. Additionally, neural–symbolic methods like Logic Tensor Networks [58] and DeepProbLog [59] offer frameworks for integrating logical constraints into the learning-based predictors and planners.
@@
Finally, hybrid architectures are designed to ensure safety and handle real-time demands. One common pattern involves a ‘safety net’ [64] approach, where a primary machine-learned policy is overseen by a simpler, verifiable safety controller that intervenes if unsafe actions are proposed. Other specialized systems focus on real-time adaptability and conflict resolution. For example, [68] formalizes explicit traffic rules to rule out infeasible maneuvers, and [61] combines incremental symbolic perception with reactive control synthesis to generate verified, rule–compliant driving maneuvers in dynamic environments.
@@
In summary, hybrid approaches that integrate symbolic and neural methods offer significant advantages by combining explicit rule-based reasoning with the flexibility of deep learning or by structuring systems with complementary learned and verifiable components. While these methods enhance interpretability and safety in complex driving scenarios, challenges remain in managing computational complexity, ensuring real-time performance, and effectively synchronizing heterogeneous components.
@@
II-F Diffusion and Other Learning-Based Prediction Incorporating Traffic Rules
@@ III Discussion and Future Directions
A central focus lies in emerging foundation model- and diffusion model approaches, which leverage unstructured data and generative mechanisms to encode traffic dynamics either via explicit symbolic modules or implicitly within learned representations. These methods promise to enhance prediction accuracy and planning robustness in complex scenarios, signaling a paradigm shift in autonomous systems. Our review equips researchers with methodologies to address prediction-planning challenges while balancing safety, efficiency, and goal achievement.
Future work can resolve real-time performance in dynamic environments, improve the interpretability of deep learning systems, and unify heterogeneous data streams (e.g., sensor inputs and traffic rules). The scalability and adaptability of foundation/diffusion models in diverse driving conditions require further validation. Multimodal reasoning and formal safety guarantees—particularly for handling nuanced, conflicting rules and hierarchical regulatory frameworks—will grow in importance.
In summary, by synthesizing state-of-the-art methods across different paradigms, this survey highlights the current capabilities and limitations in trajectory prediction and planning. It provides a roadmap for future research, emphasizing the importance of integrating explicit knowledge through symbolic reasoning, hybrid models, or generative approaches. These advancements are crucial for developing the next generation of autonomous driving systems that are both safe and efficient in real-world operations. By advancing knowledge representation techniques and fostering synergistic combinations of different approaches, we can pave the way for more intelligent, trustworthy, and safer autonomous vehicles capable of seamlessly navigating the complexities of real-world driving.